# Canonical Autoregressive Generation

### v1

@@ 4 Discussion and Future Work
In this section, we mention aspects of our work that we believe deserve additional consideration and discuss avenues for future work.
Methodology. Our main theoretical result reveals that, for BPE-, Unigram- and Wordpiece-based tokenizers, subsequences of canonical token sequences must also be canonical. This is somehow surprising since these (deterministic) tokenizers use fundamentally different tokenization techniques—BPE uses a rule based approach, Unigram maximizes the probability of the tokenization, and Wordpiece greedily encodes the text to minimize the number of tokens. Consequently, it would be very interesting to better understand what property a tokenizer needs to satisfy for our main theoretical result to hold. In this context, it would also be interesting to define relaxed notions of canonical tokenization applicable to stochastic tokenizers and adapt our main theoretical result to this type of tokenizers.
Under canonical sampling, we canonicalize the next-token distribution by redistributing the probability mass of tokens leading to non-canonical token sequences among the remaining tokens proportionally to their original probability mass. We have shown that, in comparison with the original next-token distribution, this particular canonicalized next-token distribution leads to a distribution of output sequences that is closer to the true distribution of token sequences. However, there may be other canonicalized next-token distribution leading to a distribution of output sequences that are even closer to the true distribution. In future work, it would be worth to compare different canonicalized next-token distributions, and investigate global strategies beyond (next-token) sampling to optimally redistribute the probability mass of non-canonical output token sequences.
Evaluation. As discussed at the end of Section 3.3, we cannot establish whether Theorem 2, which focuses on distributions of output token sequences, can be extended meaningfully to distributions of output strings. This limitation arises because the distribution of output strings inherently marginalizes across all possible tokenizations of any given string, creating a complex relationship between token- and string-level probability distributions. Consequently, the KL-divergence between the true distribution of strings and the distribution of output strings induced by the original next-token distribution remains theoretically intractable when compared to the corresponding KL-divergence computed in token space. Given this theoretical gap, it would be important to evaluate to which extent canonical sampling offers practical performance benefits across different state-of-the-art LLMs, particularly since users typically derive value from the string that the output token sequence represents, rather than the token sequence itself.
@@ 5 Conclusions
We have proposed canonical sampling, a simple and efficient sampling method based on the Gumbel-Max trick that ensures a given LLM generates canonical token sequences at each step of the autoregressive generation process underpinning its functioning. Further, we have shown that, in comparison with standard sampling, the distribution of token sequences generated using canonical sampling is provably closer to the true distribution of token sequences used during training.
@@
Acknowledgements. Gomez-Rodriguez acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 945719).
Appendix A Proof of Theorem 1
@@
Here, we define some additional notation regarding the BPE tokenization algorithm and prove Theorem 1.
@@
Let $\mathbf{s}=t_{1}\shortmid\dots\shortmid t_{|\mathbf{s}|}\in V^{*}$ be a tokenization of $\bm{\sigma}=c_{1}\shortmid\dots\shortmid c_{|\bm{\sigma}|}\in\Sigma^{*}$ obtained by applying merge sequence $M=(m_{1},\dots,m_{n})$. For any continuous token subsequence $\mathbf{s}^{\prime}$ of $\mathbf{s}$ spanning the substring $\bm{\sigma}^{\prime}=c_{u}\shortmid\dots\shortmid c_{v}$, $1\leq u<v\leq|\bm{\sigma}|$, the operator $[M]_{\mathbf{s}^{\prime}}$ denotes the subsequence of merges in $M$ such that $m=(r,i,j)\in[M]_{\mathbf{s}^{\prime}}$ if $m\in M$ and $u\leq i<j\leq v$.
@@
Before we prove Theorem 1, we show that the merge sequence that creates the tokenization $\mathbf{s}=\mathbf{s}_{1}\shortmid\dots\shortmid\mathbf{s}_{n}$ from a string $\bm{\sigma}$, can be partitioned into $n$ disjoint (non-continuous) subsequences of merges, that create the tokenizations $\mathbf{s}_{1},\dots,\mathbf{s}_{n}$ from the corresponding substring of $\bm{\sigma}$.
@@
Let $\mathbf{s}\in V^{*}$ be a tokenization of $\bm{\sigma}=c_{1}\shortmid\dots\shortmid c_{|\bm{\sigma}|}\in\Sigma^{*}$ obtained by applying merge sequence $M_{\mathbf{s}}$. For any partition $\mathbf{s}=\mathbf{s}_{1}\shortmid\dots\shortmid\mathbf{s}_{n}$, where $\mathbf{s}_{i}\in V^{*},i\in[n],n\in\mathbb{N}$, the following hold:
@@
    For each $\mathbf{s}_{i}\in\mathbf{s}$, there exists a merge sequence $M_{\mathbf{s}_{i}}$ such that applying $M_{\mathbf{s}_{i}}$ to the string $\texttt{dec}(\mathbf{s}_{i})$ creates $\mathbf{s}_{i}$ and $[M_{\mathbf{s}}]_{\mathbf{s}_{i}}\overset{\to}{\equiv}M_{\mathbf{s}_{i}}$.
@@
    If $\mathbf{s}_{i}$ is a tokenization of a single character $\texttt{dec}(\mathbf{s}_{i})=c$, then $[M_{\mathbf{s}}]_{\mathbf{s}_{i}}$ is the empty sequence and the statement holds trivially. Assume $\mathbf{s}_{i}$ is a tokenization of the substring $c_{u}\shortmid\dots\shortmid c_{v}$ of $\bm{\sigma}$, with $v>u>0$, and $[M_{\mathbf{s}}]_{\mathbf{s}_{i}}=(m_{1},\dots,m_{n})$. By Definition 3, all merges $m=(r,j_{1},j_{2})\in M_{\mathbf{s}}$ with $u\leq j_{1}<j_{2}\leq v$ belong in $[M_{\mathbf{s}}]_{\mathbf{s}_{i}}$, so these merges tokenize $c_{u}\shortmid\dots\shortmid c_{v}$ into $\mathbf{s}_{i}$. Then, the merge sequence $M_{\mathbf{s}_{i}}=(m^{\prime}_{1},\dots,m^{\prime}_{n})$, where for all $k\in[n]$ it holds that $m_{k}=(r,j_{1},j_{2})$ and $m^{\prime}_{k}=(r,j_{1}-u,j_{2}-u)$, $r\in\mathcal{R}$, contains the same merge rules in the same order, but with indices shifted left by $u$. So if $M_{\mathbf{s}_{i}}$ is applied to the string $\texttt{dec}(\mathbf{s}_{i})$ it will create $\mathbf{s}_{i}$.
@@
    If $[M_{\mathbf{s}}]_{\mathbf{s}_{i}}$ or $[M_{\mathbf{s}}]_{\mathbf{s}_{j}}$ are the empty sequence, meaning $\mathbf{s}_{i}$ or $\mathbf{s}_{j}$ are a tokenization of only a single character in $\bm{\sigma}$, then the statement holds trivially. If $\mathbf{s}_{i}$ is a tokenization of the substring $c_{u}\shortmid\dots\shortmid c_{v}$ and $\mathbf{s}_{j}$ is a tokenization of the substring $c_{u^{\prime}}\shortmid\dots\shortmid c_{v^{\prime}}$, since $i\neq j$ is must be that either $u<v<u^{\prime}<v^{\prime}$ or $u^{\prime}<v^{\prime}<u<v$. But for all $m=(r,i_{1},i_{2})\in[M_{\mathbf{s}}]_{\mathbf{s}_{i}}$ it holds that $u\leq i_{1}<i_{2}\leq v$, and for all $m^{\prime}=(r^{\prime},j_{1},j_{2})\in[M_{\mathbf{s}}]_{\mathbf{s}_{j}}$ it holds that $u^{\prime}\leq j_{1}<j_{2}\leq v^{\prime}$. Intuitively, it is not possible for a merge to span two different subsequences $\mathbf{s}_{i},\mathbf{s}_{j}$ in the partition of $\mathbf{s}$, because then (part of) $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$ would be merged.
@@
We can now prove Theorem 1, which we restate below.
Theorem 1. Let ${\mathcal{T}}=\{\Sigma,V,\texttt{enc},\texttt{dec}\}$ be a BPE-based tokenizer and $\mathbf{s}\in V^{*}$ a non-canonical token sequence according to ${\mathcal{T}}$. Then, for any token $t\in V$, $\mathbf{s}\shortmid t$ is also non-canonical.
@@
Assume that $\mathbf{s}\shortmid t$ is canonical. Then, there exists a unique merge sequence $M_{\mathbf{s}\shortmid t}$ that creates it following the BPE algorithm. From Lemma 1, $M_{\mathbf{s}\shortmid t}$ can be split into $[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$ and $[M_{\mathbf{s}\shortmid t}]_{t}$, where $[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$ contains the merges that create $\mathbf{s}$ and $[M_{\mathbf{s}\shortmid t}]_{t}$ contains the merges that create $t$. From Lemma 1, there exists a merge sequence $M_{\mathbf{s}}$ that creates $\mathbf{s}$ when applied to $\texttt{dec}(\mathbf{s})$ and $[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}\overset{\to}{\equiv}M_{\mathbf{s}}$. Because $\mathbf{s}$ is a prefix of $\mathbf{s}\shortmid t$, the index shift is zero and we have that $[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}=M_{\mathbf{s}}$.
Since $\mathbf{s}$ tokenized according to $[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}=(m_{1},\dots,m_{n})$ is non-canonical, there must exist a different, canonical tokenization $\mathbf{s}^{\prime}\neq\mathbf{s}$ of the same character string, $\texttt{dec}(\mathbf{s})=\texttt{dec}(\mathbf{s}^{\prime})$. Let $M_{\mathbf{s}^{\prime}}=(m^{\prime}_{1},\dots,m^{\prime}_{n^{\prime}})$ be the unique merge sequence that creates $\mathbf{s}^{\prime}$ from $\texttt{dec}(\mathbf{s})$ according to the BPE algorithm. Because $M_{\mathbf{s}^{\prime}}\neq[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$, it must be that either there exists at least one $i$, $i\leq\min(n,n^{\prime})$, such that $m_{i}\neq m^{\prime}_{i}$, or $m_{i}=m^{\prime}_{i}$ for all $i\in[\min(n,n^{\prime})]$ but $n\neq n^{\prime}$.
We will first examine the first case. Let $m_{i}\in[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$ and $m^{\prime}_{i}\in M_{\mathbf{s}^{\prime}}$ be the first merges that are different between $[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$ and $M_{\mathbf{s}^{\prime}}$, meaning $\forall j<i:\;m_{j}=m^{\prime}_{j}$, for $m_{j}\in[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$, $m^{\prime}_{j}\in M_{\mathbf{s}^{\prime}}$. Because $\mathbf{s}^{\prime}$ is canonical and $\mathbf{s}$ is not, it must be that $m^{\prime}_{i}\prec m_{i}$. We will now compare $M_{\mathbf{s}\shortmid t}$, $[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$ and $M_{\mathbf{s}^{\prime}}$. There are two sub-cases:
@@
    The first $i$ merges in $M_{\mathbf{s}\shortmid t}$ are the same as in $[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$. This means that the first $i-1$ merges are the same as in $M_{\mathbf{s}^{\prime}}$. Then, merge $m_{i}$ being applied instead of $m^{\prime}_{i}\prec m_{i}$ on substring $\texttt{dec}(\mathbf{s})$, implies that $M_{\mathbf{s}\shortmid t}$ cannot be the merge sequence that creates the canonical tokenization of $\texttt{dec}(\mathbf{s}\shortmid t)$ according to BPE.

### v2

@@ 5 Discussion and Future Work
In this section, we highlight several limitations of our work and discuss avenues for future research.
Tasks and languages. Our experiments provides strong empirical evidence that tokenization multiplicity can occur on three natural language tasks, particularly in non-english languages. However, it would be interesting to study tokenization multiplicity on additional tasks. Moreover, it would be interesting to investigate whether commonly used practices to improve multilingual language generation, such as fine-tuning on different languages, using a different tokenizer per language, or using specialized models trained on mostly non-English text, may reduce the prevalence of tokenization multiplicity.
Methodology. Our main theoretical result (Theorem 2) reveals that, for BPE-, Unigram- and Wordpiece-based tokenizers, subsequences of canonical token sequences must also be canonical. It would be very interesting to better understand what property a tokenizer needs to satisfy for our result to hold. In this context, it would also be interesting to define relaxed notions of canonical tokenization applicable to stochastic tokenizers (Kudo, 2018; Provilkov et al., 2020), and adapt our theoretical result to this type of tokenizers.
Further, under canonical generation, we canonicalize the next-token distribution by redistributing the probability mass of tokens leading to non-canonical token sequences among the remaining tokens proportionally to their original probability mass. We have shown that, in comparison with the original next-token distribution, this particular canonicalized next-token distribution leads to a distribution of output sequences that is closer to the true distribution of token sequences. However, we have found that, in practice, canonical generation has slightly lower performance than standard generation. In future work, it would be worth to investigate global strategies beyond (next-token) sampling to redistribute the probability mass of non-canonical output token sequences to achieve better practical performance.
@@ 6 Conclusions
We have presented empirical evidence that, particularly for non-english outputs, both proprietary and open-weights LLMs often generate the same (output) string with different tokenizations, even under the same input prompt, and this multiplicity of tokenizations in turn leads to arbitrary, undesirable price variation. To address the problem of tokenization multiplicity, we have proposed canonical generation, a type of constrained generation that restricts LLMs to only generate the canonical tokenization of any output string, and introduced a simple and efficient sampling algorithm based on the Gumbel-Max trick to implement it. Further, we have shown that, in comparison with standard generation, the distribution of token sequences generated using canonical generation is provably closer to the true distribution of token sequences used during training, and the performance and runtime of LLMs using either method are comparable.
@@
Acknowledgements. Gomez-Rodriguez acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 945719). Tsirtsis acknowledges support from the Alexander von Humboldt Foundation in the framework of the Alexander von Humboldt Professorship (Humboldt Professor of Technology and Regulation awarded to Sandra Wachter) endowed by the Federal Ministry of Education and Research via the Hasso Plattner Institute.
Appendix A Tokenization Algorithms
There exists many tokenization algorithms to construct the set of tokens $V$, the encoder enc, and the decoder dec characterizing a tokenizer $\mathcal{T}$}. In the following, we review three popular tokenization algorithms, BPE (Gage, 1994; Sennrich et al., 2016), Unigram (Kudo, 2018) and Wordpiece (Song et al., 2021). We also discuss pretokenization, a preprocessing technique used to partition larger bodies of text before tokenization.
A.1 The BPE tokenization algorithm
The BPE tokenization algorithm (Gage, 1994; Sennrich et al., 2016) is used by most, if not all, state-of-the-art LLMs. In a nutshell, the BPE algorithm aims to create a tokenizer $\mathcal{T}$} with a set of tokens $V$ corresponding to character sequences that appear frequently in a training set of strings $\mathcal{C}$. To this end, it proceeds as follows.
In an initialization phase, the algorithm sets i) $\Sigma$ to be the set of all characters that appear at least once in $\mathcal{C}$, ii) $V$ to be the set of single-character tokens, that is, for each $c\in\Sigma$, there exists one and only one $t\in V$ such that $\texttt{dec}(t)=c$, and iii) $\mathcal{S}$} to be the set of single-character token sequences $\mathbf{s}\in V^{+}$ representing all strings in $\mathcal{C}$. After the initialization phase, the algorithm proceeds iteratively for a predetermined number of iterations. At each iteration, it looks for the pair of tokens $t,t^{\prime}\in V$ whose concatenation $t\shortmid t^{\prime}$ appears most frequently in the set of token sequences $\mathcal{S}$}, it creates a new token $t\circ t^{\prime}$, where the symbol $\circ$ denotes the merge operation and $\texttt{dec}(t\circ t^{\prime})=\texttt{dec}(t)\shortmid\texttt{dec}(t^{\prime})$, and it adds the newly created token to $V$. Then, for each token sequence $\mathbf{s}\in{\mathcal{S}}$, it replaces all occurrences of $t\shortmid t^{\prime}$ by $t\circ t^{\prime}$ one by one. Lastly, it creates a merge rule $r_{t,t^{\prime}}$, which specifies the replacement of $t\shortmid t^{\prime}$ with $t\circ t^{\prime}$, and adds it to an ordered sequence of merge rules $\mathcal{R}$.
After termination, the algorithm defines the encoder enc and decoder dec as follows. For any given token sequence $\mathbf{s}\in V^{+}$, $\texttt{dec}(\mathbf{s})$ transforms the sequence to a string one token at a time, in order, using the token definitions. For any given string $\bm{\sigma}\in\Sigma^{+}$, $\texttt{enc}(\bm{\sigma})$ first transforms the string to a sequence of single-character tokens. Then, it merges consecutive tokens from this sequence following the merge rules from $\mathcal{R}$, in order, until no merge rule is applicable, and it returns the resulting sequence—the canonical sequence. [^13] If $t\shortmid t^{\prime}$ appears multiple times in a token sequence, the merge rule $r_{t,t^{\prime}}$ is applied in order of appearance in the sequence.
A.2 The Wordpiece tokenization algorithm
In the initialization phase, $\Sigma$ is set to contain all characters that appear at least once in the training set of strings $\mathcal{C}$. Then, for each character $c\in\Sigma$ that appears at least once in $\mathcal{C}$, a single-character token $t$ is added to $V$ such that $\texttt{dec}(t)=c$, and $\mathcal{S}$} is initialized as a set that contains all single-character token sequences $\mathbf{s}\in V^{+}$ that represent all strings in $\mathcal{C}$. Interestingly, Wordpiece transforms characters (and substrings) inside words differently than characters (and substrings) at the beginning of words. Specifically, tokens representing characters (and substrings) inside words have a special prefix.
To build the vocabulary, Wordpiece proceeds iteratively by merging existing tokens and adding them to $V$ until it reaches a predetermined size, similarly to BPE. However, the criterion to select which pair of tokens to merge is different. If $freq(\mathbf{s}^{\prime})$ denotes the number of times that sequence $\mathbf{s}^{\prime}\in V^{+}$ appears (as a subsequence) in the set of sequences $\mathcal{S}$}, Wordpiece looks for the pair of tokens $t,t^{\prime}\in V$ that maximizes the value of $\frac{freq(t\shortmid t^{\prime})}{freq(t)\cdot freq(t^{\prime})}$. Then, a new token $t\circ t^{\prime}$ is added to $V$, where $\texttt{dec}(t\circ t^{\prime})=\texttt{dec}(t)\shortmid\texttt{dec}(t^{\prime})$, and all occurences of $t\shortmid t^{\prime}$ in each token sequence $\mathbf{s}\in{\mathcal{S}}$ are replaced by $t\circ t^{\prime}$. With this criterion, Wordpiece prefers to merge tokens whose concatenation appears commonly in $\mathcal{S}$}, but they are not common individually.
After the above iterative process terminates, the algorithm defines the encoder and decoder functions as follows. For any token sequence $\mathbf{s}=t_{1}\shortmid\dots\shortmid t_{n}\in V^{+}$ with $n\in\mathbb{N}$, the decoder returns $\texttt{dec}(\mathbf{s})=\texttt{dec}(t_{1})\shortmid\dots\shortmid\texttt{dec}(t_{n})$ using the token definitions. Any string $\bm{\sigma}=c_{1}\shortmid\dots\shortmid c_{m}\in\Sigma^{+}$, $m\in\mathbb{N}$ given to the encoder is tokenized greedily from left to right, each time selecting the token in the vocabulary that represents the most characters starting from the beginning of the string. Specifically, the first token in $\texttt{enc}(\bm{\sigma})$ is the token $t\in V$ such that $\texttt{dec}(t)=c_{1}\shortmid\dots\shortmid c_{i}$, with $i\leq m$, and $\nexists t^{\prime}\in V$ such that $\texttt{dec}(t^{\prime})=c_{1}\shortmid\dots\shortmid c_{j}$ with $i<j\leq m$. In the above selection, if $c_{1}$ is inside a word, then $t$ must contain the special prefix. This process continues in the same manner with the remaining string $c_{i+1}\shortmid\dots\shortmid c_{m}$.
A.3 The Unigram tokenization algorithm
After the initial vocabulary has been set, the algorithm proceeds in iterations, each time computing a loss over the strings in $\mathcal{C}$ and the current vocabulary, and removing a batch of tokens from the vocabulary (typically $10\%$ or $20\%$ of tokens) whose removal minimizes this loss. In each iteration, every token $t$ in the current vocabulary $V$ is assigned a probability score $r(t)=\frac{freq(t)}{\sum_{t^{\prime}\in V}freq(t^{\prime})}$, where $freq(t)$ denotes the number of times that the token $t$ appears in all possible tokenizations of the strings in $\mathcal{C}$. For each token $t\in V$, the loss over the training set is computed as $\sum_{\bm{\sigma}\in\mathcal{C}}-\log(r_{V\setminus\{t\}}(\bm{\sigma}))$, where $r_{V}(\bm{\sigma})=\max_{\mathbf{s}\in V^{+},\texttt{dec}(\mathbf{s})=\bm{\sigma}}r(\mathbf{s})$ denotes the probability score of the most likely tokenization of $\bm{\sigma}$ under vocabulary $V$, and the probability score of tokenization $\mathbf{s}=t_{1}\shortmid\dots\shortmid t_{n}$, with $n\in\mathbb{N}$, is simply $r(\mathbf{s})=r(t_{1})\dots r(t_{n})$. The tokens that minimize this loss are removed from the vocabulary and the process repeats until the vocabulary reaches a predetermined size.
After the vocabulary has been finalized, the encoder is set to tokenize a string $\bm{\sigma}\in\Sigma^{+}$ by finding its most likely tokenization under the final vocabulary $V$, i.e., $\texttt{enc}(\bm{\sigma})=\arg\max_{\mathbf{s}\in V^{+},\texttt{dec}(\mathbf{s})=\bm{\sigma}}r(\mathbf{s})$, using the Viterbi algorithm (Viterbi, 1967), and the decoder decodes all tokens in $V$ the same way as in the original, large vocabulary.
Appendix B Proof of Theorem 2
In this section, we prove Theorem 2 by showing individually that each tokenization algorithm—BPE, Wordpiece, and Unigram—builds tokenizers which are non-recovering.
B.1 BPE-based tokenizers are non-recovering
@@
In order to show that BPE-based tokenizers are non-recovering, we define some additional notation regarding the BPE tokenization algorithm.
@@
Let $\mathbf{s}=t_{1}\shortmid\dots\shortmid t_{|\mathbf{s}|}\in V^{+}$ be a tokenization of $\bm{\sigma}=c_{1}\shortmid\dots\shortmid c_{|\bm{\sigma}|}\in\Sigma^{+}$ obtained by applying merge sequence $M=(m_{1},\dots,m_{n})$. For any continuous token subsequence $\mathbf{s}^{\prime}$ of $\mathbf{s}$ spanning the substring $\bm{\sigma}^{\prime}=c_{u}\shortmid\dots\shortmid c_{v}$, $1\leq u<v\leq|\bm{\sigma}|$, the operator $M]_{\mathbf{s}^{\prime}}$ denotes the subsequence of merges in $M$ such that $m=(r,i,j)\in[M]_{\mathbf{s}^{\prime}}$ if $m\in M$ and $u\leq i<j\leq v$.
@@
Before we prove that BPE-based tokenizers are non-recovering, we show that the merge sequence that creates the tokenization $\mathbf{s}=\mathbf{s}_{1}\shortmid\dots\shortmid\mathbf{s}_{n}$ from a string $\bm{\sigma}$, can be partitioned into $n$ disjoint (non-continuous) subsequences of merges, that create the tokenizations $\mathbf{s}_{1},\dots,\mathbf{s}_{n}$ from the corresponding substring of $\bm{\sigma}$.
@@
Let $\mathbf{s}\in V^{+}$ be a tokenization of $\bm{\sigma}=c_{1}\shortmid\dots\shortmid c_{|\bm{\sigma}|}\in\Sigma^{+}$ obtained by applying merge sequence $M_{\mathbf{s}}$. For any partition $\mathbf{s}=\mathbf{s}_{1}\shortmid\dots\shortmid\mathbf{s}_{n}$, where $\mathbf{s}_{i}\in V^{+},i\in[n],n\in\mathbb{N}$, the following hold:
@@
    For each $\mathbf{s}_{i}\in\mathbf{s}$, there exists a merge sequence $M_{\mathbf{s}_{i}}$ such that applying $M_{\mathbf{s}_{i}}$ to the string $\texttt{dec}(\mathbf{s}_{i})$ creates $\mathbf{s}_{i}$ and $M_{\mathbf{s}}]_{\mathbf{s}_{i}}\overset{\to}{\equiv}M_{\mathbf{s}_{i}}$.
@@
    If $\mathbf{s}_{i}$ is a tokenization of a single character $\texttt{dec}(\mathbf{s}_{i})=c$, then $M_{\mathbf{s}}]_{\mathbf{s}_{i}}$ is the empty sequence and the statement holds trivially. Assume $\mathbf{s}_{i}$ is a tokenization of the substring $c_{u}\shortmid\dots\shortmid c_{v}$ of $\bm{\sigma}$, with $v>u>0$, and $M_{\mathbf{s}}]_{\mathbf{s}_{i}}=(m_{1},\dots,m_{n})$. By Definition 4, all merges $m=(r,j_{1},j_{2})\in M_{\mathbf{s}}$ with $u\leq j_{1}<j_{2}\leq v$ belong in $M_{\mathbf{s}}]_{\mathbf{s}_{i}}$, so these merges tokenize $c_{u}\shortmid\dots\shortmid c_{v}$ into $\mathbf{s}_{i}$. Then, the merge sequence $M_{\mathbf{s}_{i}}=(m^{\prime}_{1},\dots,m^{\prime}_{n})$, where for all $k\in[n]$ it holds that $m_{k}=(r,j_{1},j_{2})$ and $m^{\prime}_{k}=(r,j_{1}-u,j_{2}-u)$, $r\in\mathcal{R}$, contains the same merge rules in the same order, but with indices shifted left by $u$. So if $M_{\mathbf{s}_{i}}$ is applied to the string $\texttt{dec}(\mathbf{s}_{i})$ it will create $\mathbf{s}_{i}$.
@@
    If $M_{\mathbf{s}}]_{\mathbf{s}_{i}}$ or $M_{\mathbf{s}}]_{\mathbf{s}_{j}}$ are the empty sequence, meaning $\mathbf{s}_{i}$ or $\mathbf{s}_{j}$ are a tokenization of only a single character in $\bm{\sigma}$, then the statement holds trivially. If $\mathbf{s}_{i}$ is a tokenization of the substring $c_{u}\shortmid\dots\shortmid c_{v}$ and $\mathbf{s}_{j}$ is a tokenization of the substring $c_{u^{\prime}}\shortmid\dots\shortmid c_{v^{\prime}}$, since $i\neq j$ it must be that either $u<v<u^{\prime}<v^{\prime}$ or $u^{\prime}<v^{\prime}<u<v$. But for all $m=(r,i_{1},i_{2})\in[M_{\mathbf{s}}]_{\mathbf{s}_{i}}$ it holds that $u\leq i_{1}<i_{2}\leq v$, and for all $m^{\prime}=(r^{\prime},j_{1},j_{2})\in[M_{\mathbf{s}}]_{\mathbf{s}_{j}}$ it holds that $u^{\prime}\leq j_{1}<j_{2}\leq v^{\prime}$. Intuitively, it is not possible for a merge to span two different subsequences $\mathbf{s}_{i},\mathbf{s}_{j}$ in the partition of $\mathbf{s}$, because then (part of) $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$ would be merged.
@@
Building up on Lemma 1, we show that BPE-based tokenizers are non-recovering.
Lemma 2.
BPE-based tokenizers are non-recovering.
@@
Assume that $\mathbf{s}\shortmid t$ is canonical. Then, there exists a unique merge sequence $M_{\mathbf{s}\shortmid t}$ that creates it following the BPE algorithm. From Lemma 1, $M_{\mathbf{s}\shortmid t}$ can be split into $M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$ and $M_{\mathbf{s}\shortmid t}]_{t}$, where $M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$ contains the merges that create $\mathbf{s}$ and $M_{\mathbf{s}\shortmid t}]_{t}$ contains the merges that create $t$. From Lemma 1, there exists a merge sequence $M_{\mathbf{s}}$ that creates $\mathbf{s}$ when applied to $\texttt{dec}(\mathbf{s})$ and $M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}\overset{\to}{\equiv}M_{\mathbf{s}}$. Because $\mathbf{s}$ is a prefix of $\mathbf{s}\shortmid t$, the index shift is zero and we have that $M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}=M_{\mathbf{s}}$.
Since $\mathbf{s}$ tokenized according to $M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}=(m_{1},\dots,m_{n})$ is non-canonical, there must exist a different, canonical tokenization $\mathbf{s}^{\prime}\neq\mathbf{s}$ of the same character string, $\texttt{dec}(\mathbf{s})=\texttt{dec}(\mathbf{s}^{\prime})$. Let $M_{\mathbf{s}^{\prime}}=(m^{\prime}_{1},\dots,m^{\prime}_{n^{\prime}})$ be the unique merge sequence that creates $\mathbf{s}^{\prime}$ from $\texttt{dec}(\mathbf{s})$ according to the BPE algorithm. Because $M_{\mathbf{s}^{\prime}}\neq[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$, it must be that either there exists at least one $i$, $i\leq\min(n,n^{\prime})$, such that $m_{i}\neq m^{\prime}_{i}$, or $m_{i}=m^{\prime}_{i}$ for all $i\in[\min(n,n^{\prime})]$ but $n\neq n^{\prime}$.
We will first examine the first case. Let $m_{i}\in[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$ and $m^{\prime}_{i}\in M_{\mathbf{s}^{\prime}}$ be the first merges that are different between $M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$ and $M_{\mathbf{s}^{\prime}}$, meaning $\forall j<i:\;m_{j}=m^{\prime}_{j}$, for $m_{j}\in[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$, $m^{\prime}_{j}\in M_{\mathbf{s}^{\prime}}$. Because $\mathbf{s}^{\prime}$ is canonical and $\mathbf{s}$ is not, it must be that $m^{\prime}_{i}\prec m_{i}$. We will now compare $M_{\mathbf{s}\shortmid t}$, $M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$ and $M_{\mathbf{s}^{\prime}}$. There are two sub-cases:
@@
    The first $i$ merges in $M_{\mathbf{s}\shortmid t}$ are the same as in $M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$. This means that the first i1i-1 merges are the same as in $M_{\mathbf{s}^{\prime}}$. Then, merge $m_{i}$ being applied instead of $m^{\prime}_{i}\prec m_{i}$ on substring $\texttt{dec}(\mathbf{s})$, implies that $M_{\mathbf{s}\shortmid t}$ cannot be the merge sequence that creates the canonical tokenization of $\texttt{dec}(\mathbf{s}\shortmid t)$ according to BPE.